OpenStudy (anonymous):

A ball is dropped from a cliff that is 224 feet high. The distance S (in feet) that it falls in t seconds in given by the formula S=224-16tsquare. How many seconds (to tenths) will it take for the ball to hit the ground?

OpenStudy (hunus):

I think S should be the height of the ball after t seconds, because at t=0 'S' is 224. So you want to find out when the height of the ball h is equal to zero h = 224 - 16t^2 0 = 224 - 16t^2

OpenStudy (hunus):

You there, Dria?